MINUTES FROM THE BOOSTER SUPERVISOR MEETING
DATE: June 25, 2002
PSB home page, PSB
PRESENT: A. Blas, C. Carli, G. Cyvoct, M. O’Neil, R. Scrivens
Supervisor Linac: G. GrŠwer ˆ M. O’Neil
Supervisor PSB: E. Jensen ˆ C. Carli
- Sunday : Due to a thunderstorm,
the HT of the source and the RF Hazemeyer went off and the LI vacuum
- Monday : Increased flashover
rate of the source. This coincides with the arrival of wet weather as observed
already previously. (Note that at the moment, due to dust created by the
installation of the Laser for LIS, the air there does not come from the
old Faraday cage, where humidity is controlled, but from the tunnel.) caused
by air from the tunnel, which is not humidity controlled, laser
installation creates some dust).
- General remark on steering in
the Linac transfer lines :
- Steering at the exit
of tank 3 of Linac 2 is very delicate. Due to the observation of significant
“hot spots”, the trajectory has been adjusted to minimize radiation
monitor signals and thus activation.
- LT steering dipoles 10
to 30 should only be adjusted by Linac supervisors
- Steering now at
elevated radiation levels, to be redone next week in MD time.
- Radiation alarm due to timing
problems of the extraction kickers. Caused by PS-DSC’s in fault leading to
an erroneous frequency sent to the PSB for synchronization.
- Again problems with the
active filter of the extraction septum power supply.
- A control electronics card of
the BT.BHZ10 power supply has been exchanged (as mentioned in the minutes
from the last meeting), because the CCV and AQN values were rather
different. As a consequence,
the CCV values had to be readjusted (356 A → 361 A for
destination PS and –375 A → -378 A for destination BDUMP or
ISOLDE) and a slow drift has been observed during the week.
- Vistar problems are
- High intensity in Ring 1 : Losses
during the first 50 ms significantly higher than normal. Re-optimisation
by M. Chanel of the resonance 3QV=16 allowed to improve. Furthermore
a problem with the transverse damper was observed and further
investigations should be done as soon as possible.
- The baseline shift of analog
signals of the pick-ups BT.UES30 and BT.UES40 with high intensity (a high
intensity bunch causes a sudden jump of the baseline, eventually at the
passage of another bunch) has been observed again.
- Bad injection efficiency on
all users observed. Simultaneously and independently, the DSC DPSBINRG was
rebooted and a faulty TG8 module in the same DSC was exchanged. It is not
clear which action resolved the problem.
- Note that, if the DSC
DPSBVACU is off, all other DSC’s of the PSB cannot be rebooted from the
- The next BS meeting will take
place on Tuesday July 2, 2002 at 10.00 in the MCR Glassbox.
Minutes by C. Carli